### 27 Apr 2026 v2

# Radial evolution of Alfvén wave Parametric Decay Instability in the near-Sun solar wind: Effects of Temperature Anisotropy

Abstract
Parametric decay instability (PDI) of Alfvén wave is thought to play an important role in the dissipation of the large-amplitude Alfvén waves and in the heating of magnetized plasmas. Temperature anisotropy is frequently observed by spacecrafts, including Parker Solar Probe (PSP), in the near-Sun solar wind, yet its impact on PDI in the near-Sun solar wind has been understudied. We calculate the maximum growth rates of PDI, $\gamma_{\max}/\omega_{0}$, where $\omega_{0}$ is the frequency of the parent wave, by solving the linear dispersion relation of Chew-Goldberger-Low (CGL) equations under several expanding background models. To assess the effect of temperature anisotropy, the growth rate is compared with that derived from ideal magnetohydrodynamics (MHD). From $R_{0}$ ($1.02R_{\odot}$) to $30R_{0}$, we consider three expansion cases: (i) spherically symmetric adiabatic expansion with constant wind speed, (ii) Multi-source observation- and model-constrained expansion, and (iii) a PSP-constrained profile of $(\beta_{\parallel},\xi)$, where $\beta_{\parallel}=8\pi p_{\parallel 0}/B_{0}^{2}$ is the parallel plasma beta and $\xi=T_{\perp 0}/T_{\parallel 0}$ is the temperature anisotropy, that includes Parker-spiral effects. We find that temperature anisotropy increases $\gamma_{\max}/\omega_{0}$ for $\beta\lesssim 0.1$ in the near-Sun solar wind: in the case of (iii), temperature anisotropy with $T_{\perp 0}>T_{\parallel 0}$ increases $\gamma_{\max}/\omega_{0}$ by factors of $\sim 1.5$ over $R\simeq 1$$10\,R_{0}$, whereas temperature anisotropy with $T_{\parallel 0}>T_{\perp 0}$ decreases $\gamma_{\max}/\omega_{0}$ at larger $R$. Our results suggest that the temperature anisotropy plays an important role in the onset of PDI even in low-$\beta$ regimes, such as the near-Sun solar wind.
show]saguchi.hayato.s4@dc.tohoku.ac.jp
hide]kawazura@a.utsunomiya-u.ac.jp
hide]shoda.m.astroph@gmail.com
hide]yuto.katoh@tohoku.ac.jp
I Introduction
One of the major unresolved problems in solar physics is how the solar corona is heated and how the solar wind is accelerated. It is widely believed that the energy source of these phenomena comes from energy of magneto-convection on the photosphere, and Alfvén waves carry the energy to heat the solar corona and drive the solar wind. In fact, some studies suggest that Alfvén waves observed in the chromosphere and low corona have enough energy to drive the solar wind (De Pontieu et al., 2007; McIntosh et al., 2011). The mechanism of how Alfvén waves propagate and dissipate is, however, still under discussion.
Reflection-driven Alfvén wave turbulence has been extensively studied as a candidate process for coronal heating and the solar wind acceleration by Alfvén waves (e.g., Matthaeus et al., 1999; Dmitruk et al., 2002; Cranmer et al., 2007; Verdini and Velli, 2007; Perez and Chandran, 2013; van der Holst et al., 2014; van Ballegooijen and Asgari-Targhi, 2016, 2017; Usmanov et al., 2018; Chandran et al., 2025a; Meyrand et al., 2025). Due to the gradients of the background Alfvén speed in the solar corona and the solar wind, Alfvén waves are partially reflected (Heinemann and Olbert, 1980; Velli, 1993). The nonlinear interaction between sunward and anti-sunward propagating Alfvén waves occurs, leading to the onset of Alfvén wave turbulence, which drives energy to small scales, ultimately heating the plasma. (Howes and Nielson, 2013). Previous studies suggest that Alfvén wave turbulence model based on reduced MHD model may be insufficient to provide enough energy to sustain the solar wind (e.g., Perez and Chandran, 2013), whereas small-scale gradients due to density fluctuations can increase the heating rate (van Ballegooijen and Asgari-Targhi, 2016, 2017). Indeed, reduced MHD model that incorporates observationally constrained density fluctuations showed sufficient Alfvén wave reflection and turbulence to heat coronal holes and drive the fast solar wind (Asgari-Targhi et al., 2021). From this perspective, several studies have included compressibility into Alfvén wave turbulence model and have succeeded in reproducing coronal heating and the fast solar wind self-consistently (Shoda and Yokoyama, 2018; Shoda et al., 2018a, b, 2019; Matsumoto, 2021).
As a physical process for generating the density fluctuation, enhancing reflection rate and promoting Alfvénic turbulence (e.g., Suzuki and Inutsuka, 2006; Shoda et al., 2019; Matsumoto, 2021), parametric decay instability (PDI) of Alfvén wave has attracted attention. In PDI, a large-amplitude forward propagating Alfvén wave (referred to as mother wave, parent wave, or pump wave) decays into daughter waves: a backward propagating Alfvén wave and a forward propagating slow-magnetoacoustic wave (Galeev and Oraevskii, 1963; Sagdeev and Galeev, 1969). Goldstein (1978) and Derby (1978) derived the dispersion relation of PDI from ideal MHD equations. The solar corona and the near-Sun solar wind are a favorable environment for the development of PDI because the instability preferentially occurs in low-$\beta$ plasmas.
PDI has been studied extensively over the past several decades including dispersive effects (e.g., Wong and Goldstein, 1986; Longtin and Sonnerup, 1986; Hollweg, 1994), kinetic effects (e.g., Inhester, 1990; Araneda, 1998; Nariyuki and Hada, 2006, 2007), two-fluid (e.g., Viñas and Goldstein, 1991b, a), expansion effects (e.g., Matteini et al., 2010; Tenerani and Velli, 2013), the suppressing effects of turbulence on PDI (e.g., Shi et al., 2017; Fu et al., 2018), temperature anisotropy (e.g., Hamabata, 1993; Tenerani et al., 2017), relativistic effects (Ishizaki and Ioka, 2024), and the parametric decay of kinetic Alfvén waves (e.g., Hasegawa and Chen, 1976; Li et al., 2024). PDI has also been investigated in a variety of settings and with diverse numerical approaches, including studies of nonlinear evolution (e.g., Terasawa et al., 1986; Hoshino and Goldstein, 1989), non-monochromatic waves (e.g., Umeki and Terasawa, 1992; Malara and Velli, 1996), multi-dimensional numerical simulations (e.g., Del Zanna et al., 2001, 2015), fully kinetic particle-in-cell (PIC) simulations (e.g., Nariyuki et al., 2008; González et al., 2023), and Alfvénic wave packets (e.g., Tenerani et al., 2020; Marriott and Tenerani, 2024a, b; Komissarov, 2025). In addition, several studies have examined secondary effects triggered by PDI, such as inverse cascade (e.g., Réville et al., 2018; Chandran, 2018), anisotropic heating via stochastic heating (Comişel et al., 2019), and PDI-driven turbulence (e.g., Shoda and Yokoyama, 2018; Shoda et al., 2019). Several studies have attempted to identify observational evidence of PDI. The signatures of PDI were reported in the solar transition region (Hahn et al., 2022), and several other works have presented suggestive results (Bowen et al., 2018; Kasper et al., 2021; Hahn et al., 2025). Although Li et al. (2025) proposed a novel and experimentally viable scheme to quantify the growth of PDI recently, observational examples of PDI remain relatively scarce (Zank et al., 2022; Zhao et al., 2022).
Temperature anisotropy is frequently observed by Parker Solar Probe (PSP) in the near-Sun solar wind (e.g., Verniero et al. 2020; Woodham et al. 2021; Huang et al. 2020, 2023a, 2023b, 2025; Short et al. 2024; Laker et al. 2024; Yogesh et al. 2025). In some intervals, the observed anisotropy becomes high, with $T_{\perp}/T_{\parallel}\gtrsim 10$ around the ion cyclotron instability threshold, where $\perp$ and $\parallel$ denote the perpendicular and parallel components to the magnetic field (e.g., Coello-Guzmán et al., 2026). Theoretical modeling of species-dependent temperatures and temperature anisotropies in the solar wind remains an active and important topic (e.g., Ofman et al., 2001, 2002; Meng et al., 2015; Chandran et al., 2025b; Yoon et al., 2026). In particular, several mechanisms have been proposed to account for them, such as the helicity barrier (e.g., Meyrand et al., 2021; Squire et al., 2022; McIntyre et al., 2025; Panchal et al., 2025) and the stochastic heating, both of which have attracted considerable attention as an efficient channel for perpendicular heating in the solar wind (e.g., Chandran et al., 2010, 2013; Cerri et al., 2021; Bowen et al., 2025; Johnston et al., 2025; Mallet et al., 2026). In global solar-wind modeling, Alfvén Wave Solar Model (AWSoM; van der Holst et al. 2014; Meng et al. 2015; van der Holst et al. 2022) has achieved notable success in predicting temperature anisotropies via stochastic heating.
Regarding the relationship between temperature anisotropy and PDI, however, theoretical studies that explicitly quantify its effects on PDI are scarce. The pioneering analyses of Hamabata (1993) and Tenerani et al. (2017) employ linear theory based on the Chew-Goldberger-Low (CGL) equations (Chew et al., 1956) and show that a background temperature anisotropy enhances the growth rate when $T_{\perp}>T_{\parallel}$ and reduces it when $T_{\perp}<T_{\parallel}$. Nevertheless, these works do not address the regime most relevant to the near-Sun solar wind—namely, low-$\beta$ conditions in which PSP observes large temperature anisotropies—leaving its impact on the growth rate largely unexplored. Furthermore, the heliocentric locations where PDI is expected to occur have been investigated by computing the radial evolution of the maximum growth rate in expanding isotropic solar-wind models (e.g., Umeki and Terasawa, 1992; Tenerani and Velli, 2013; Del Zanna et al., 2015; Réville et al., 2018; Shoda et al., 2018b, 2019). However, the radial evolution of the maximum growth rate based on PSP observations remains poorly explored, and, to date, no study has incorporated temperature anisotropy within such a framework. As a result, our understanding of how the pronounced near-Sun anisotropy influences the onset and radial evolution of PDI remains unclear.
In this study, following Tenerani et al. (2017), we analyze the maximum growth rate of PDI in the near-Sun solar wind, incorporating temperature anisotropy. Specifically, we address the following questions.
  1. 1.
    How much does the temperature anisotropy enhance the maximum growth rate of PDI in the low-$\beta$ region?
  2. 2.
    In the adiabatic expansion, does the temperature anisotropy affect the radial evolution of the maximum growth rate of PDI?
  3. 3.
    In the expansion cases based on the observed data of PSP, does the temperature anisotropy affect the radial evolution of the maximum growth rate of PDI?
Additionally, we discuss the effect of the perturbed temperature anisotropy on the maximum growth rate of PDI.
II method
II.1 Governing equations and dispersion relations
We use two dispersion relations of the parametric instability that are derived respectively from their basic equations, the isotropic one (ideal MHD) and the anisotropic one (CGL equations, Chew et al., 1956) to compare two results and evaluate the effect of temperature anisotropy. Dispersion relations can be derived by linearizing the governing equations about a finite-amplitude mother wave and retaining small perturbations corresponding to the daughter waves and the background plasma (Goldstein, 1978; Derby, 1978). The governing equations in the isotropic case (ideal MHD) are as follows:
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})=0,$ (II.1)
$\rho\left(\frac{\partial}{\partial t}+\mathbf{u}\cdot\nabla\right)\mathbf{u}=-\nabla p+\frac{(\nabla\times\mathbf{B})\times\mathbf{B}}{4\pi},$ (II.2)
$\frac{\partial\mathbf{B}}{\partial t}=\nabla\times(\mathbf{u}\times\mathbf{B}),$ (II.3)
$\frac{dp}{d\rho}=c_{s}^{2}.$ (II.4)
$p\rho^{-n}=\rm{const},$ (II.5)
where the parameters in a set of equations are as follows:
  • $\rho$ denotes the mass density.
  • $\mathbf{u}$ denotes the vector of the velocity.
  • $p$ denotes the pressure.
  • $\mathbf{B}$ denotes the vector of the magnetic field.
  • $c_{s}$ denotes the sound speed.
  • $n$ denotes the polytropic index. In this study, we assume adiabaticity, so we set $n$ as $5/3$.
After linearization, the dispersion relation is as follows (Goldstein, 1978; Derby, 1978):
[…]
where the parameters in the dispersion relation are as follows:
  • $\hat{\omega}\equiv\omega/\omega_{0}+i\,\gamma/\omega_{0}$,
    the (dimensionless) daughter-wave complex frequency, normalized by the mother-wave frequency $\omega_{0}$.
  • $\hat{k}\equiv k/k_{0}$,
    the (dimensionless) daughter-wave wavenumber, normalized by the mother-wave wavenumber $k_{0}$.
  • $\beta\equiv 8\pi p_{0}/B_{0}^{2}$,
    the plasma beta, defined as the ratio of the gas pressure to the magnetic pressure.
  • $v_{A}\equiv B_{0}/\sqrt{4\pi\rho_{0}}$,
    the Alfvén speed defined in terms of the background magnetic-field strength $B_{0}$ and mass density $\rho_{0}$.
  • $\hat{B}_{\perp}\equiv B_{\perp}/B_{0}$,
    the mother-wave transverse magnetic-field amplitude $B_{\perp}$ normalized by the background field $B_{0}$.
In the anisotropic case, we adopt the Chew–Goldberger–Low (CGL) equations, which describe temperature anisotropy in a weakly collisional plasma under double-adiabatic conditions (Chew et al., 1956). CGL equations are as follows:
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})=0,$ (II.7)
$\begin{split}\rho\left(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}\right)&=-\nabla\left(p_{\perp}+\frac{B^{2}}{8\pi}\right)\\ &\quad+\frac{1}{4\pi}\,\mathbf{B}\cdot\nabla\mathbf{B}+\nabla\cdot\!\left(\hat{\mathbf{b}}\hat{\mathbf{b}}\,\Delta p\right),\end{split}$ (II.8)
$\frac{\partial\mathbf{B}}{\partial t}=\nabla\times(\mathbf{u}\times\mathbf{B}),$ (II.9)
$\left(\frac{\partial}{\partial t}+\mathbf{u}\cdot\nabla\right)\left(\frac{p_{\parallel}B^{2}}{\rho^{3}}\right)=0,$ (II.10)
$\left(\frac{\partial}{\partial t}+\mathbf{u}\cdot\nabla\right)\left(\frac{p_{\perp}}{\rho B}\right)=0,$ (II.11)
where the new parameters in a set of equations are as follows:
  • $p_{\perp}$ denotes the pressure perpendicular to the magnetic field.
  • $p_{\parallel}$ denotes the pressure parallel to the magnetic field.
  • $\mathbf{\hat{b}}=\textbf{B}/B$ denotes the unit vector along the magnetic field.
  • $\Delta p=p_{\perp}-p_{\parallel}$ denotes the pressure anisotropy, the difference between $p_{\perp}$ and $p_{\parallel}$.
After linearization, the dispersion relation derived from CGL equations (Tenerani et al., 2017) is as follows:
$\begin{split}\left[\hat{\omega}^{2}-\tilde{\beta}\hat{k}^{2}\left(1+\frac{\hat{B}_{\perp}^{2}\xi}{3}\right)\right]\left\{(\hat{\omega}-\hat{k})\left[(\hat{\omega}+\hat{k})^{2}-4\right]+\frac{\tilde{\beta}\hat{B}_{\perp}^{2}(\xi-4)}{3\left(1+\hat{B}_{\perp}^{2}\right)}\left[(\hat{k}^{2}+1)\hat{\omega}+\hat{k}(\hat{k}^{2}-3)\right]\right\}\\ =\hat{B}_{\perp}^{2}\hat{k}^{2}\left[1-\frac{\tilde{\beta}\left(3-\xi-\hat{B}_{\perp}^{2}\right)}{3\left(1+\hat{B}_{\perp}^{2}\right)}\right]\left\{\hat{\omega}^{3}+\hat{\omega}^{2}\hat{k}-3\hat{\omega}+\hat{k}-\frac{\tilde{\beta}(3-\xi)}{3}\left[(\hat{k}^{2}+1)\hat{\omega}+\hat{k}(\hat{k}^{2}-3)\right]\right\},\end{split}$ (II.12)
where the plasma parameters in the dispersion relation are as follows:
  • $\hat{\omega},\hat{k}$ and $\hat{B}_{\perp}$ are the same as those of the ideal-MHD case.
  • $\beta_{\parallel}=8\pi p_{\parallel 0}/B_{0}^{2}$,
    the parallel plasma beta.
  • $\xi=T_{0\perp}/T_{\parallel 0}$,
    the background temperature anisotropy.
  • $\tilde{\beta}=3\beta_{\parallel}/[2(1+\hat{B}_{\perp}^{2}+\beta_{\parallel}(\xi-1)/2)]$.
Note that we do not solve the full set of governing equations numerically, but solve two dispersion relations by prescribing the background model beforehand in this study.
Table 1: Background solar wind models
Evolution of $B_{0}$ and $\rho_{0}$ Evolution of $T_{\perp 0}$ and $T_{\parallel 0}$ Evolution of $\hat{B}_{\perp}^{2}$
Case 1 (Spherically symmetric adiabatic expansion with constant wind speed) $B_{0}\propto R^{-2}$ $\rho_{0}\propto R^{-2}$ Anisotropic: $\frac{d}{dt}(p_{\parallel 0}B_{0}^{2}/\rho_{0}^{3})=0$ & $\frac{d}{dt}(p_{\perp 0}/\rho_{0}B_{0})=0$ $\rightarrow\ T_{\parallel 0}=const,\ T_{\perp 0}\propto R^{-2}$ Isotropic: $T_{0}\rho_{0}^{-2/3}=const\rightarrow\ T_{0}\propto R^{-4/3}$ $\hat{B}_{\perp}^{2}\propto R^{1}$
Case 2 (Multi-source observation- and model-constrained expansion) Profiles derived from the analysis of Metis and PSP observations (Telloni et al. (2021)) Anisotropic: The result of AWSoM (Meng et al. (2015)) Isotropic: $T_{0}=(2T_{\perp 0}+T_{\parallel 0})/3$ with the anisotropic profile $\hat{B}_{\perp}^{2}\propto R^{0.96}$ (Huang et al. (2023c))
Case 3 (PSP-constrained expansion) Profiles derived from PSP observations (Short et al. (2024)) Anisotropic: Profiles derived from PSP observations (Short et al. (2024)) Isotropic: $T_{0}=(2T_{\perp 0}+T_{\parallel 0})/3$ with the anisotropic profile $\hat{B}_{\perp}^{2}\propto R^{0.96}$ (Huang et al. (2023c))
II.2 Background solar wind models
As plasma emanating from the Sun flows radially outward, it undergoes expansion, and the background plasma parameters evolve with the radial distance. To investigate the radial evolution of $\gamma_{\rm{max}}/\omega_{0}$ under various solar-wind conditions, we set three radial background models in the range from $R_{0}$ ($1.02R_{\odot}$) to $30R_{0}$. In each case, we prescribe the magnetic field $B_{0}$, mass density $\rho_{0}$, temperature (either $T_{0}$ or $T_{\parallel 0}$ and $T_{\perp 0}$), and $\hat{B}_{\perp}$ as functions of the normalized radial distance $\hat{R}=R/R_{0}$ (where $R_{0}=1.02R_{\odot}$ denotes the lower coronal boundary), based on theoretical assumptions and, where appropriate, results from previous studies. We assume that $\omega_{0}$ does not depend on $R$ although some previous studies consider it (e.g., Tenerani and Velli, 2013). In other words, the present setup does not explicitly incorporate the effect of solar-wind acceleration. We also assume that these parameters depend only on $R$. Table 1 shows three expansion cases considered in this study.
II.2.1 Case 1: Spherically symmetric adiabatic expansion with constant wind speed
In case 1, both the isotropic and anisotropic models assume adiabatic evolution; we examine how the difference between an MHD-type adiabatic assumption and a CGL-type (double-adiabatic) assumption affects the radial evolution of $\gamma_{\rm{max}}/\omega_{0}$.
In terms of the background magnetic field $B_{0}$ and the background mass density $\rho_{0}$ in case 1, we assume spherical expansion and a steady solar wind ($\partial/\partial t=0$, ${v}_{r}=\mathrm{const}$). Then $\nabla\!\cdot\!\mathbf{B}=0$ and mass conservation equation lead to
$B_{0}\propto R^{-2},\ \rho_{0}\propto R^{-2},$ (II.13)
respectively (Grappin et al., 1993; Grappin and Velli, 1996; Matteini et al., 2024). This applies to both the isotropic and anisotropic formulations.
Table 2: Boundary values at the corona
$\beta(R_{0})$ $\xi(R_{0})$ $\hat{B}_{\perp}^{2}(R_{0})$
Case 1 (Spherically symmetric adiabatic expansion with constant wind speed) 0.001, 0.01, 0.1 10 0.01, 0.001
Case 2 (Multi-source observation- and model-constrained expansion) Determined by Telloni et al. (2021) and Meng et al. (2015) Determined by Meng et al. (2015) 0.01, 0.001
Case 3 (PSP-constrained expansion) Determined by Short et al. (2024) Determined by Short et al. (2024) 0.01, 0.001
In terms of the temperature anisotropy, we apply the double-adiabatic closures such as equations of (II.10) and (II.11) to the background. Then we get
$\frac{T_{\parallel 0}B_{0}^{2}}{\rho_{0}^{2}}=\mathrm{const},\quad\frac{T_{\perp 0}}{B_{0}}=\mathrm{const}.$ (II.14)
By combining this with $B_{0}\propto R^{-2},\ \rho_{0}\propto R^{-2}$, we can obtain the radial evolution of $T_{\parallel 0}$ and $T_{\perp 0}$, such as:
$T_{\parallel 0}=\mathrm{const},\quad T_{\perp 0}\propto R^{-2}.$ (II.15)
In the isotropic case, we assume $n=5/3$ and use the adiabatic equation of state
$T_{0}\,\rho_{0}^{-2/3}=\mathrm{const},$ (II.16)
to obtain the radial evolution of the isotropic temperature. By combining this with $\rho_{0}\propto R^{-2}$, the radial evolution of $T_{0}$ is
$T_{0}\propto R^{-4/3}.$ (II.17)
In terms of the mother wave amplitude $B_{\perp}$, we adopt
$B_{\perp}\propto R^{-1.5},$ (II.18)
based on Wentzel–Kramers–Brillouin (WKB) theory (e.g., Whang, 1973). By combining this with $B_{0}\propto R^{-2}$ we can obtain
$\hat{B}_{\perp}^{2}\propto R.$ (II.19)
These radial evolutions follow those expected from the Expanding Box Model (EBM) (Grappin et al., 1993; Grappin and Velli, 1996; Matteini et al., 2024). Although observations indicate that the solar wind undergoes non-adiabatic expansion (Marsch et al., 1982, 1983; Hellinger et al., 2011, 2013), several numerical studies have investigated PDI while partly assuming adiabatic expansion for investigating the effects of expansion on the wave propagation and the solar wind dynamics (e.g., Tenerani and Velli, 2013; Réville et al., 2018). In the expanding box model (EBM), it is well known that in the low-$\beta$ regime, the maximum growth rate scales with heliocentric distance as $\gamma_{\rm{\max}}/\omega_{0}\propto R^{1/3}$ (e.g., Tenerani and Velli, 2013; Réville et al., 2018). Although adiabatic expansion is the idealized expansion case, analyzing the radial evolution of PDI under this simplified assumption provides a useful benchmark and enables comparison with previous work.
II.2.2 Case 2: Multi-source observation- and model-constrained expansion
[…]
$\hat{B}_{\perp}^{2}\propto R^{0.96}.$ (II.20)
This radial evolution is almost equal to the equation (II.19), which is also consistent with results from AWSoM simulation (van der Holst et al., 2022). To our knowledge, no observationally derived radial profile of temperature anisotropy within $\sim 10R_{\odot}$ is currently available; therefore, we use the AWSoM simulation results reported by Meng et al. (2015). The goal of this case is to quantify how strongly temperature anisotropy can influence PDI under an observation- and model- constrained solar-wind background, but note that this case does not include the $B_{\theta}$ and $B_{\phi}$ components in the background magnetic field.
II.2.3 Case 3: PSP-constrained expansion
In case 3, we adopt (i) the PSP-based radial profile of $(\beta_{\parallel},\,\xi)$ based on Short et al. (2024) and (ii) a radial profile of the magnetic-fluctuation amplitude inferred from PSP measurements by Huang et al. (2023c). The goal of this scenario is also to quantify how strongly temperature anisotropy can influence PDI under the more observationally constrained solar-wind background, which includes the Parker spiral in the background magnetic field. In Appendix A, we show the detail of the profiles in this case.
II.3 Boundary values
To prescribe the radial profiles of the background parameters, we need to impose boundary conditions at the lower boundary. Table 2 shows the imposed boundary values at the lower boundary in all cases. For case 1, we adopt three values of the parallel plasma beta, $\beta(R_{0})=10^{-1},\,10^{-2},\,10^{-3}$ (Gary, 2001). On the other hand, for case 2 and 3, the coronal value of the parallel plasma beta is determined by Telloni et al. (2021); Meng et al. (2015) or Short et al. (2024). For case 1, we set the temperature anisotropy to $\xi(R_{0})\equiv T_{\perp 0}/T_{\parallel 0}=10$ (Telloni et al., 2007; Cranmer et al., 2008; Meng et al., 2015; Cranmer, 2020), and for case 2 or 3, the temperature anisotropy is determined by Meng et al. (2015) or Short et al. (2024). In terms of the coronal values of the normalized mother wave amplitude, we adopt $\hat{B}_{\perp}^{\,2}(R_{0})=10^{-2}$ in the corona (Réville et al., 2018; Hahn et al., 2022; van der Holst et al., 2022; Huang et al., 2023c).
III Results
Figure 1: Radial evolution of the maximum growth rate and the corresponding background plasma parameters for case 1 with $\beta(R=R_{0})=0.001,\,0.01,\,0.1$, $\xi(R_{0})=10$, and $\hat{B}_{\perp}^{2}(R_{0})=0.01$. The upper panel shows the ratio of the maximum growth rate in the CGL model to that in the ideal MHD model, $\gamma_{\max}^{\mathrm{CGL}}/\gamma_{\max}^{\mathrm{MHD}}$. Three gray horizontal dashed lines indicate $0.5,1,1.5$ respectively. The solid, dashed, and dotted lines correspond to $\beta(R_{0})=0.001$, $0.01$, and $0.1$, respectively. The middle panel presents the radial evolution of the maximum growth rate in both models: the green curves denote the ideal-MHD results, whereas the orange curves denote the CGL results. The solid black line is an auxiliary line corresponding to $\gamma_{\max}/\omega_{0}\propto R^{1/4}$. Green and orange solid, dashed, and dotted lines correspond to $\beta(R_{0})=0.001$, $0.01$, and $0.1$, respectively. The lower panel shows the radial evolution of the local background parameters used to evaluate the maximum growth rate. The purple curve indicates the temperature anisotropy $\xi\equiv T_{\perp 0}/T_{\parallel 0}$ (here $\xi(R_{0})=10$), and the black curve indicates $\hat{B}_{\perp}^{2}$ (here $\hat{B}_{\perp}^{2}(R_{0})=0.01$); these two profiles are common to both the ideal-MHD and CGL models. The red and blue curves represent $\beta_{\parallel}$ and $\beta$, respectively, and the solid/dashed/dotted line styles follow the same convention as in the middle panel.
Figure 2: Radial evolution of the maximum growth rate and the corresponding background plasma parameters for case 1 with $\beta(R_{0})=0.001,\,0.01,\,0.1,\,$ $\xi(R_{0})=10$, and $\hat{B}_{\perp}^{2}(R_{0})=0.001$. The figure format is identical to that of Figure 1.
Figures 1 and 2 show the radial evolution of the PDI maximum growth rate in case 1 for two values of the mother-wave amplitude, $\hat{B}_{\perp}^{2}(R_{0})=0.01$ and $0.001$, respectively. In each figure, the upper panel shows the ratio $\gamma_{\max}^{\mathrm{CGL}}/\gamma_{\max}^{\mathrm{MHD}}$, the middle panel shows the radial evolution of $\gamma_{\max}/\omega_{0}$ in the two models, and the lower panel shows the corresponding radial profiles of the background parameters used in the calculation.
From the results shown in the upper panels of both Figures 1 and 2, we find that, in the near-Sun region ($R<2R_{0}$), temperature anisotropy with $T_{\perp}>T_{\parallel}$ increases the PDI maximum growth rate, whereas at larger distances ($R<2R_{0}$) temperature anisotropy with $T_{\parallel}>T_{\perp}$ suppresses it. This trend is consistently confirmed for all three beta cases ($\beta(R_{0})=0.1,\,0.01,\,0.001$) and for both amplitude cases ($\hat{B}_{\perp}^{2}(R_{0})=0.01,\,0.001$), and it becomes more pronounced as $\beta$ increases. Because the absolute value of the maximum growth rate is smaller at higher $\beta$, the effect of temperature anisotropy appears relatively larger.
For the MHD results, the radial evolution of the maximum growth rate in the middle panels of Figures 1 and 2 shows that, for the two low-$\beta$ cases ($\beta(R_{0})=0.001$ and $0.01$), the maximum growth rate tends to increase with $R$. In the lowest-$\beta$ case of $\hat{B}_{\perp}^{2}(R_{0})=0.01$, the growth rate approximately follows the black solid guideline $\propto R^{1/4}$. Previous studies (e.g., Jayanti and Hollweg, 1993) estimated, based on analytical expressions, that under spherically symmetric adiabatic expansion, the maximum growth rate scales as $\propto R^{1/3}$, whereas our results exhibit a slightly different scaling. One plausible reason for this discrepancy is that $\hat{B}_{\perp}^{2}$ is not sufficiently small for the small-amplitude assumption to be strictly valid. Indeed, the smaller-amplitude case (Figure 2) is closer to the $R^{1/3}$ scaling than the case of Figure 1.
Regarding comparisons with previous studies that examined the radial evolution of the maximum growth rate, the MHD results under a radial evolution of EBM are broadly consistent with earlier works adopting similar assumptions. The tendency for the maximum growth rate to increase with $R$ as a power law in the low-$\beta$ regime agrees with the results of Réville et al. (2018). The weak decrease of the maximum growth rate with $R$ for the $\beta(R_{0})=0.1$ case is consistent with the results of Tenerani and Velli (2013), who explicitly included the $R$-dependence of the pump frequency, as well as with those of Del Zanna et al. (2015), who did not include it (noting, however, that Del Zanna et al. (2015) used a somewhat different value, $\hat{B}_{\perp}^{2}=0.04$ at the inner boundary).
For the CGL results, the middle panels of Figures 1 and 2 indicate that, for all three beta cases, the maximum growth rate tends to decrease with $R$. This implies that, within an EBM-CGL framework, the maximum growth rate decays with heliocentric distance, in contrast to the adiabatic EBM-MHD case where it tends to increase with $R$.
Figure 3: Radial evolution of the maximum growth rate and the background plasma parameters for case 2 with $\hat{B}_{\perp}^{2}(R_{0})=0.01$ and $0.001$. The upper panel shows the ratio of the maximum growth rate in the CGL model to that in the ideal MHD model, $\gamma_{\max}^{\mathrm{CGL}}/\gamma_{\max}^{\mathrm{MHD}}$. Three gray horizontal dashed lines indicate $0.5,1,1.5$ respectively. The solid and dashed lines correspond to $\hat{B}_{\perp}^{2}(R_{0})=0.01$ and $0.001$, respectively. The middle panel presents the radial evolution of the maximum growth rate in both models: the green curves denote the ideal-MHD results, whereas the orange curves denote the CGL results. The lower panel shows the radial evolution of the local background parameters used to evaluate the maximum growth rate. The purple curve indicates the temperature anisotropy $\xi\equiv T_{\perp 0}/T_{\parallel 0}$. The black curve indicates $\hat{B}_{\perp}^{2}$ (here $\hat{B}_{\perp}^{2}(R_{0})=0.01,0.001$); these two profiles are common to both the ideal-MHD and CGL models. The red and blue curves represent $\beta_{\parallel}$ and $\beta$, respectively.
Figure 4: Radial evolution of the maximum growth rate and the background plasma parameters for case 3 with $\hat{B}_{\perp}^{2}(R_{0})=0.01$ and $0.001$ in the common profile of $\beta$ and $T_{\perp 0}/T_{\parallel 0}$. The figure format is identical to that of Figure 3.
Figure 3 shows the results for case 2 with $\hat{B}_{\perp}^{2}(R_{0})=0.01$ and $0.001$. In the upper panel, the ratio of the maximum growth rates reaches $\sim 1.5$ at $R\sim 2R_{0}$, and then increases up to $\sim 4$ for $\hat{B}_{\perp}^{2}(R_{0})=0.01$ and up to $\sim 20$ for $\hat{B}_{\perp}^{2}(R_{0})=0.001$. Thus, from the upper panel of Figure 3, we find that a temperature anisotropy with $T_{\perp}>T_{\parallel}$ enhances the maximum growth rate in both cases, $\hat{B}_{\perp}^{2}(R_{0})=0.001$ and $0.01$. From the middle panel, we find that the PDI is unstable only within $R\lesssim 10R_{0}$. As indicated by the lower panel, this confinement is caused by the rapid increase of $\beta$ and the rapid decrease of $\xi$ associated with solar-wind expansion.
Figure 4 shows the results for case 3, i.e., the PSP-constrained expansion, again in the same three-panel format as Figure 3. For the upper panel of Figure 4, we confirm again that temperature anisotropy with $T_{\perp}>T_{\parallel}$ increases the maximum growth rate by approximately a factor of $\sim 1.5$ within $R<10R_{0}$. This result demonstrates that the temperature-anisotropy-driven enhancement of the maximum growth rate remains effective even in a low-$\beta$ regime with $\beta_{\parallel}\lesssim 0.1$. Also, it is clear that, at larger $R$, the maximum PDI growth rate is reduced due to the suppression effects caused by a parallel-dominant temperature anisotropy and by the perturbed temperature anisotropy, which is discussd in Section IV.2.
Previous studies investigate the radial evolution of the PDI maximum growth rate under conditions closer to the non-adiabatic actual solar wind (Shoda et al., 2018b, 2019). The results of cases 2 and 3 show a faster decrease of the maximum growth rate with increasing $R$ than in Shoda et al. (2018b, 2019); however, the qualitative feature that the growth rate attains a local maximum—as reported in both Shoda et al. (2018b) and Shoda et al. (2019)—is reproduced. We attribute the faster decline of $\gamma_{\max}$ with increasing $R$ in case 2 to a more rapid increase in $\beta$ than that found in both Shoda et al. (2018b) and Shoda et al. (2019). A plausible reason for this difference is that those studies considered polar fast wind, which is expected to be less dense and therefore to remain at lower $\beta$ than the low-latitude wind represented by case 2.
IV Discussion
IV.1 Parameter scans
To clarify the physical origin of the $\beta_{\parallel}$-dependence and $\xi$-dependence of the PDI growth rate, we present three parameter-scan figures. Figure 5 compares $\gamma_{\max}/\omega_{0}$ in ideal MHD and CGL under the same background conditions, with $\xi=1$ and $\hat{B}_{\perp}=0.1$ fixed in both models, so that the horizontal axis can be written as the common plasma beta, $\beta(=\beta_{\parallel})$. Figure 6 shows a complementary comparison for an anisotropic background: in the CGL model, $\xi=10$ and $\hat{B}_{\perp}=0.1$ are fixed and $\gamma_{\max}/\omega_{0}$ is plotted against $\beta_{\parallel}$, whereas in ideal MHD, $\hat{B}_{\perp}=0.1$ is fixed and $\gamma_{\max}/\omega_{0}$ is plotted against $\beta$. Figure 7 isolates the direct $\xi$-dependence of the CGL growth rate by plotting $\gamma_{\max}^{\mathrm{CGL}}/\omega_{0}$ as a function of $\xi$ at fixed $\beta_{\parallel}=0.01$ and $\hat{B}_{\perp}=0.1$.
In Figure 5, we find that in the low-$\beta$ regime ($\beta\lesssim 10^{-2}$) $\gamma_{\max}/\omega_{0}$ is nearly identical between MHD and CGL. In contrast, for $\beta\gtrsim 10^{-2}$, the MHD maximum growth rate becomes systematically larger than the CGL value. This difference reflects the effect of perturbed temperature anisotropy, which is discussed in Section IV.2.
Figure 6 illustrates how a temperature anisotropy with $T_{\perp}>T_{\parallel}$ modifies the overall $\beta$-dependence of the PDI growth rate in CGL relative to that in ideal-MHD. As shown in Figure 6, the two models are again similar at sufficiently low $\beta$, whereas toward the higher-$\beta$ side of the scan, the PDI growth rate in CGL remains systematically larger than the ideal-MHD value when $\xi=10$ is imposed. The results shown in Figures 5 and 6 indicate that, in the low-$\beta$ regime, the dependence of the maximum growth rate on $\beta$ in ideal MHD is essentially identical to its dependence on $\beta_{\parallel}$ in CGL, even when $\xi=10$ is imposed.
Figure 7 shows the direct $\xi$-dependence of the PDI growth rate in CGL at fixed $\beta_{\parallel}=0.01$ and $\hat{B}_{\perp}=0.1$. As shown in this figure, $\gamma_{\max}^{\mathrm{CGL}}/\omega_{0}$ varies only weakly over a broad range of $\xi$, and a more pronounced increase appears only at sufficiently large $\xi$.
From these results, we interpret the case 1 behavior as follows. In case 1, different adiabatic assumptions were adopted for MHD and CGL, which led to distinct radial scalings: $\beta\propto R^{2/3}$ in MHD, whereas $\beta_{\parallel}\propto R^{2}$ in CGL. The increasing difference between $\beta$ and $\beta_{\parallel}$ with $R$ explains the opposite radial behavior of $\gamma_{\max}$ in MHD and CGL.
Figure 5: Maximum growth rate in MHD (black dots) and in CGL (red dots) as a function of $\beta$, evaluated for $\hat{B}_{\perp}=0.1$ and $\xi=1$.
Figure 6: Maximum growth rate versus plasma beta in ideal MHD (black dots) and versus parallel plasma beta in CGL with $\xi=10$ (red dots), both for $\hat{B}_{\perp}=0.1$.
Figure 7: Maximum growth rate in CGL, $\gamma_{max}^{\rm{CGL}}/\omega_{0}$ (red dots), as a function of $\xi$, evaluated for $\hat{B}_{\perp}=0.1$ and $\beta_{\parallel}=0.01$.
IV.2 The effect of the perturbed temperature anisotropy
Figure 8: This figure presents a contour plot of $\gamma_{\max}^{\rm CGL}/\gamma_{\max}^{\rm MHD}$ (contour interval $\Delta=0.1$) for $\xi=1$. The vertical axis shows $\hat{B}_{\perp}$ over the range $[0.01,1]$, and the horizontal axis shows $\beta$ over the range $10^{-4},10^{2}]$.
The differences between the maximum growth rates obtained from the isotropic and anisotropic dispersion relations arise from both the background temperature anisotropy and the anisotropy of the perturbations ($\delta T_{\perp}\neq\delta T_{\parallel}$). Even if we impose a background temperature anisotropy of unity in the CGL-based dispersion relation, (II.12), the resulting dispersion relation remains completely different from the isotropic (ideal-MHD) dispersion relation, (II.6). The reason why the dispersion relations differ even under identical background conditions is that the pressure perturbations—i.e., the linearized pressure tensor $\delta{P}$—are not actually the same in the ideal-MHD and CGL models. In other words, because the two linearized sets of governing equations are not equivalent, the dispersion relations derived from them are also different. However, it can be shown that the two linearized systems become identical if, in addition to setting $\xi=1$, we further impose $\delta T_{\parallel}=\delta T_{\perp}$, provided that the polytropic index on the MHD side is assumed to be $5/3$ (Appendix B).
Therefore, allowing $\delta T_{\parallel}$ and $\delta T_{\perp}$ to differ introduces the difference between the two dispersion relations under the same background. To the best of our knowledge, this study is the first to quantitatively demonstrate the impact of perturbed temperature anisotropy by taking the difference between the ideal-MHD and CGL dispersion relations of the parametric instability.
[…]
IV.3 Limitations of this study
Our study is based on solving linear dispersion relations and therefore does not provide a fully self-consistent description of the coupled evolution of the background plasma and the waves. We quantify how prescribed background properties modify the wave dynamics, whereas the reciprocal feedback of the waves on the background is not captured. Consequently, dedicated simulations that evolve both directions of the coupling are required to assess whether, and to what extent, temperature anisotropy regulates PDI.
Since PDI can enhance the generation of counter-propagating Alfvén waves, it is expected to reduce cross helicity. Since temperature anisotropy with $T_{\perp}>T_{\parallel}$ promotes the PDI and increases the generation of counter-propagating waves, the resulting reduction in cross helicity may weaken the helicity barrier, thereby suppressing anisotropic heating. In addition, the relationship between PDI and stochastic heating has been discussed in the literature (Comişel et al., 2019). At present, it remains unclear whether signatures of temperature-anisotropy-enhanced PDI can be identified in the actual solar wind. Further validation from both numerical simulations and in-situ observations is therefore required.
A variety of physical processes not included here can also modify the PDI maximum growth rate, including kinetic effects (e.g., collisionless damping such as Landau damping), large-scale expansion and acceleration, and suppression by background turbulence. In reality, these effects operate simultaneously, making it difficult to quantify the net effect of the PDI. Because these processes act to reduce the PDI growth rate, it is plausible that the effective region where PDI operates is more restricted than the range inferred from our case 3. Moreover, these effects have often been studied in isolation; a natural next step is to assess PDI in models that integrate multiple physical processes in a unified framework.
The expansion cases considered in this work are somewhat idealized, and there remains room to construct more unideal, observationally constrained expansion cases. In particular, the radial evolution of $\xi$ in the innermost heliosphere is not well captured owing to limited observational coverage. We anticipate that improved empirical constraints on the radial evolution of temperature anisotropy will become feasible as the availability of in-situ measurements continues to increase. For example, observationally based radial trends classified by solar-wind speed and heliographic latitude, and ideally also by source region, would provide valuable constraints.
V Conclusion and Summary
We investigated how the temperature anisotropy modifies the onset and radial evolution of the Alfvén wave parametric decay instability (PDI) in the near-Sun solar wind. We prescribed expanding background profiles and computed the normalized maximum growth rate, $\gamma_{\max}/\omega_{0}$, by directly solving two linear dispersion relations: (i) the ideal-MHD dispersion relation of Goldstein (1978); Derby (1978) and (ii) the CGL-based dispersion relation of Tenerani et al. (2017). To consider the role of expansion, we considered multiple expansion cases over $1\leq R/R_{0}\leq 30$ and emphasized three representative cases in the main text: case 1 (spherically symmetric adiabatic expansion with constant wind speed), case 2 (Multi-source observation- and model-constrained expansion) and case 3 (PSP-constrained expansion).
Our main findings are summarized as follows.
  1. 1.
    Background temperature anisotropy can significantly enhance PDI even in low-$\beta$ near-Sun conditions. In case 3, where $\beta_{\parallel}\lesssim 0.1$ in the inner heliosphere, the CGL model yields a maximum growth rate larger than the ideal-MHD value by a factor of $\sim 1.5$ over $R/R_{0}\sim 1$$10$ when the background anisotropy is perpendicular-temperature dominated. Moreover, in case 3, temperature anisotropy with $T_{\parallel}>T_{\perp}$ reduces the maximum growth rate, thereby limiting the radial extent of the PDI-unstable region to $R\lesssim 30R_{0}$. Other effects—such as solar-wind expansion, background turbulence, and Landau damping—are not considered in this study, so our estimated radial range for PDI occurrence could be confined to even smaller heliocentric distances. Our result demonstrates that anisotropy-driven enhancement remains effective in the parameter regime most relevant to PSP observations, and therefore temperature anisotropy should be treated as a non-negligible control parameter when assessing where PDI can operate in the near-Sun wind.
  2. 2.
    In the adiabatic case, the radial evolution of $\gamma_{\max}/\omega_{0}$ differs qualitatively between the isotropic and anisotropic closures. Under the adiabatic assumption, ideal MHD predicts that $\gamma_{\max}/\omega_{0}$ increases with heliocentric distance in the low-$\beta$ regime, with an approximate power-law scaling that approaches the classical ${\gamma}_{\max}/\omega_{0}\propto R^{1/3}$ behavior in the smaller-amplitude limit. In contrast, the CGL model does not exhibit a monotonic increase with $R$. Instead, ${\gamma}_{\max}/\omega_{0}$ can turn over and decrease beyond a certain distance. The parameter scans in Section IV.1 suggest that in the low-$\beta$ regime, the PDI growth rates in both models are nearly identical even when $\xi=10$ is imposed, whereas the direct $\xi$-dependence itself remains weak over a broad range. Because the adiabatic closures imply different radial scalings, $\beta\propto R^{2/3}$ in MHD and $\beta_{\parallel}\propto R^{2}$ in CGL, the more rapid increase of $\beta_{\parallel}$ in the CGL closure provides a plausible explanation for the different radial behaviors found in case 1.
  3. 3.
    Perturbed temperature anisotropy reduces ${\gamma}_{\max}/\omega_{0}$. Even when the background anisotropy is set to unity ($\xi=1$), the CGL-based dispersion relation does not reduce to the ideal-MHD form because the linearized pressure responses differ unless one further imposes $\delta T_{\perp}=\delta T_{\parallel}$ (for $n=5/3$). By comparing ${\gamma}_{\max}/\omega_{0}$ from CGL at $\xi=1$ with the ideal-MHD result, we quantified the impact of the perturbed temperature anisotropy. The resulting maps show that ${\gamma}_{\max}/\omega_{0}$ is generally smaller in CGL than in MHD, and the difference grows with increasing $\hat{B}_{\perp}$ and $\beta$. This implies that allowing $\delta T_{\perp}\neq\delta T_{\parallel}$ tends to suppress PDI.
These results have direct implications for PDI-based interpretations and models of the near-Sun solar wind. Because PDI produces compressive fluctuations and counter-propagating Alfvénic components, any temperature-anisotropy-driven modulation of ${\gamma}_{\max}/\omega_{0}$ can influence the generation of density fluctuations, the effective reflection rate, and the subsequent development of Alfvén-wave turbulence. Our findings therefore motivate incorporating anisotropic thermodynamics when estimating PDI activity in inner-heliosphere applications, including global simulations that use PDI as an effective mechanism for controlling turbulence.
Finally, several caveats and future directions follow naturally from this work. Our calculations are based on linear dispersion relations with prescribed backgrounds; thus, the two-way feedback between wave evolution (including PDI-driven compressive and backward components) and the background thermodynamics is not captured. Moreover, additional processes known to reduce PDI growth in the solar wind—such as expansion and acceleration, turbulence, dispersive/Hall physics, and kinetic damping—were not included and may further confine the effective PDI operating region closer to the Sun than suggested by our case 3. A key next step is to test temperature-anisotropy-modified PDI using simulations that evolve both the wave dynamics and the anisotropic background self-consistently.
This work was financially supported by the WISE Program for Sustainability in the Dynamic Earth (SyDE), Tohoku University. The work of YK is supported by JSPS KAKENHI grant 25K00975. MS is supported by JSPS KAKENHI Grant Number JP24K00688. HS thanks K. Isono, K. Saito and W. Ishizaki for the supports on code development.
{contribution}
H. Saguchi led the analysis and manuscript preparation. Y. Kawazura, M. Shoda, and Y. Katoh provided valuable discussions and feedback and reviewed the entire manuscript.
Appendix A Profiles that we use: Telloni et al. 2021, Meng et al. 2015 and Short et al. 2024
[…]
$\log_{10}B(\hat{R})=-0.901205(\log_{10}\hat{R})^{4}+3.566625(\log_{10}\hat{R})^{3}-3.448904(\log_{10}\hat{R})^{2}-2.331466(\log_{10}\hat{R})+6.003316.$ (A1)
[…]
$\log_{10}n(\hat{R})=2.8999(\log_{10}\hat{R})^{4}-11.2618(\log_{10}\hat{R})^{3}+16.7644(\log_{10}\hat{R})^{2}-13.3988(\log_{10}\hat{R})+8.9892.$ (A2)
[…]
where
$L_{\rm rise}(\hat{R})=\left[1+\exp\left\{-k_{1}\left(\ln\hat{R}-\mu_{1}\right)\right\}\right]^{-1},$ (A4)
$L_{\rm fall}(R)=\left[1+\exp\left\{-k_{2}\left(\ln R-\mu_{2}\right)\right\}\right]^{-1}.$ (A5)
Here, $A$ represents the background anisotropy level, $D$ controls the amplitude of the enhancement, and $(k_{1},\mu_{1})$ and $(k_{2},\mu_{2})$ set the steepness and the characteristic locations of the rise and fall, respectively. The parameters were obtained by a bounded nonlinear least-squares fit to the extracted data points, yielding the best-fit values
$(A,\,D,\,k_{1},\,\mu_{1},\,k_{2},\,\mu_{2})=(1.445318,\,50.000000,\,16.603014,\,0.710795,\,1.930799,\,0.321134).$ (A6)
To construct a smooth radial profile of the parallel proton temperature, $T_{\parallel}(R)$, we digitized the data points from Meng et al. (2015) within $R_{0}$$30R_{0}$ and fitted them with a smoothing spline in logarithmic radius. Specifically, we first sorted the extracted pairs $\{(R_{i},\,T_{\parallel,i})\}$ in ascending $R$ and introduced the logarithmic coordinate
$x\equiv\ln\!\left(\frac{R}{R_{\odot}}\right).$ (A7)
We then modeled $T_{\parallel}$ as a smooth function of $x$ using a cubic smoothing spline,
$T_{\parallel}(\hat{R})=S(x),$ (A8)
where $S(x)$ is obtained by minimizing the standard smoothing-spline functional
$\sum_{i=1}^{N}\left[T_{\parallel,i}-S(x_{i})\right]^{2}\;+\;s\int\left[S^{\prime\prime}(x)\right]^{2}\,dx,$ (A9)
[…]
$\beta_{\parallel}(\hat{R})=10^{f_{\beta}(\hat{R})},\qquad f_{\beta}(r)=a\hat{R}^{3}+b\hat{R}^{2}+c\hat{R}+d,$ (A10)
where the coefficients are
$a$ $=-8.71798336\times 10^{-5},$
$b$ = $=\ \ 2.68031491\times 10^{-3},$
$c$ = $=\ \ 1.01522939\times 10^{-1},$
$d$ $=-3.10411607.$ (A11)
This construction guarantees $\beta_{\parallel}(\hat{R})>0$ for all $\hat{R}$ and reproduces the anchor values $\beta_{\parallel}(1)=10^{-3}$, $\beta_{\parallel}(13)=3\times 10^{-2}$, $\beta_{\parallel}(20)=2\times 10^{-1}$, and $\beta_{\parallel}(30)=1$.
[…]
$\xi(\hat{R})=\xi_{1}\exp\!\left[-I(\hat{R})\right],\qquad\xi_{1}\equiv\xi(1)=15,$ (A12)
with
$\frac{dI}{d\hat{R}}=\left(u\hat{R}^{2}+v\hat{R}+w\right)^{2}\geq 0.$ (A13)
[…]
$I(\hat{R})=A_{5}\left(\hat{R}^{5}-1\right)+A_{4}\left(\hat{R}^{4}-1\right)+A_{3}\left(\hat{R}^{3}-1\right)+A_{2}\left(\hat{R}^{2}-1\right)+A_{1}\left(\hat{R}-1\right),$ (A14)
so that
$\xi(\hat{R})=15\exp\!\left\{-\left[A_{5}(\hat{R}^{5}-1)+A_{4}(\hat{R}^{4}-1)+A_{3}(\hat{R}^{3}-1)+A_{2}(\hat{R}^{2}-1)+A_{1}(\hat{R}-1)\right]\right\}.$ (A15)
The coefficients used in this work are
$A_{5}$ $=8.503658213699558\times 10^{-7},$
$A_{4}$ $=-5.682864240998027\times 10^{-5},$
$A_{3}$ $=8.835153670285457\times 10^{-4},$
$A_{2}$ $=5.1814868111597565\times 10^{-3},$
$A_{1}$ $=8.836698570459859\times 10^{-3}.$ (A16)
These coefficients reproduce the anchor values $\xi(1)=15$, $\xi(13)=3$, $\xi(20)=0.8$, and $\xi(30)=0.5$ exactly.
Appendix B Origin of the difference between the ideal MHD and CGL dispersion relations at $\xi=1$
First, we set only the background temperature anisotropy to unity and show that, even in this case, the linearized pressure tensor is not identical. The pressure tensor in CGL is
$\boldsymbol{P}=p_{\perp}\boldsymbol{I}+(p_{\parallel}-p_{\perp})\boldsymbol{b}\boldsymbol{b}.$ (B1)
Here, $\mathbf{I}$ is the identity tensor and $\boldsymbol{b}=\boldsymbol{B}/B$ is the unit vector along the local magnetic-field direction. We decompose the variables into a background and small perturbations as
$\rho=\rho_{0}+\delta{\rho},\quad p_{\perp}=p_{\perp 0}+\delta p_{\perp},\quad p_{\parallel}=p_{\parallel 0}+\delta p_{\parallel},\quad\boldsymbol{b}=\boldsymbol{b}_{0}+\delta\boldsymbol{b},$ (B2)
where $\boldsymbol{b_{0}}$ includes the background magnetic field $B_{0}$ and the mother wave amplitude $B_{\perp}$. In this case, the zeroth-order pressure tensor is
$\boldsymbol{P}_{0}=p_{\perp 0}\boldsymbol{I}+(p_{\parallel 0}-p_{\perp 0})\boldsymbol{b}_{0}\boldsymbol{b}_{0}.$ (B3)
The first-order pressure tensor is
$\delta{\boldsymbol{P}}=\delta{p_{\perp}}\boldsymbol{I}+(\delta{p_{\parallel}}-\delta{p_{\perp}})\boldsymbol{b}_{0}\boldsymbol{b}_{0}+(p_{\parallel 0}-p_{\perp 0})(\boldsymbol{b}_{0}\delta{\boldsymbol{b}}+\delta{\boldsymbol{b}\boldsymbol{b}_{0}}).$ (B4)
On the other hand, when the linearization in ideal MHD is assumed to be
$\rho=\rho_{0}+\delta{\rho},\quad p=p_{0}+\delta{p},\quad\boldsymbol{b}=\boldsymbol{b_{0}}+\delta{\boldsymbol{b}}.$ (B5)
The pressure tensor in ideal MHD is
$\boldsymbol{P}=p\boldsymbol{I}.$ (B6)
So, the zeroth-order pressure tensor is
$\boldsymbol{P}_{0}=p_{0}\boldsymbol{I},$ (B7)
and the first-order pressure tensor is
$\delta{\boldsymbol{P}}=\delta{p}\boldsymbol{I}.$ (B8)
Therefore, even though we assume $\xi=p_{\perp 0}/p_{\parallel 0}$ = 1, the pressure in CGL becomes
$\delta{\boldsymbol{P}}=\delta{p_{\perp}}\boldsymbol{I}+(\delta{p_{\parallel}}-\delta{p_{\perp}})\boldsymbol{b}_{0}\boldsymbol{b}_{0}+\cancel{(p_{\parallel 0}-p_{\perp 0})(\boldsymbol{b}_{0}\delta{\boldsymbol{b}}+\delta{\boldsymbol{b}\boldsymbol{b}_{0}})},$ (B9)
not equal to the pressure tensor of ideal MHD, the equation (B8) due to $\delta{p}_{\perp}\neq\delta{p}_{\parallel}$. So, roughly speaking, the linearized set of governing equations in CGL is not the same as ideal MHD (though the differential equations for pressure in CGL are not the same as in ideal MHD). The dispersion relation of the parametric instability is derived from a set of linearized equations of governing equations, so the dispersion relation in CGL is not equal to the dispersion relation in ideal MHD even if $\xi=1$.
If we only assume $\xi=1$, three of the governing equations, such as the mass conservation, the momentum equation, and the induction equation in CGL like equations (II.7)–(II.9), are equal to three of equations in ideal MHD like equations (II.1)–(II.3). So, these three linearized equations in CGL are also equal to three linearized equations in ideal MHD when $\xi=1$.
It should also be emphasized that imposing $\xi=1$ and $\delta p_{\parallel}=\delta p_{\perp}$ is not sufficient for the linearized governing equations to become identical between the ideal-MHD and CGL frameworks. In addition, one must assume a polytropic closure with the polytropic index $n=5/3$; otherwise, the linearized equation sets do not coincide. We demonstrate this point below. Linearized equations of equations (II.10), (II.11) are
$\frac{\delta{p_{\perp}}}{p_{\perp}}=\frac{\delta{\rho}}{\rho_{0}}+\frac{\delta{B}}{B_{0}},\quad\frac{\delta{p_{\parallel}}}{p_{\parallel}}=\frac{3\delta{\rho}}{\rho_{0}}-\frac{2\delta{B}}{B_{0}}.$ (B10)
Then if we assume $p_{\perp}=p_{\parallel}=p$ and $\delta p_{\parallel}=\delta p_{\perp}=\delta{p}$, equations (B10) become like
$\frac{\delta{p}}{\delta{\rho}}=\frac{5}{3}\frac{p_{0}}{\rho_{0}}.$ (B11)
This linearized relation is equivalent to the linearized ideal-MHD closure in equation (II.4) only when the polytropic index is set to $n=5/3$.
So, if we assume the polytropic index $n=5/3$, $p_{\perp}=p_{\parallel}=p$ and $\delta p_{\parallel}=\delta p_{\perp}=\delta{p}$, a set of linearized equations in CGL is equal to a set of linearized equations in ideal MHD; therefore, the dispersion relation in CGL is equal to the dispersion relation in ideal MHD. For these reasons, even when we set $n=5/3$ and $\xi=1$, the maximum growth rates obtained from the two dispersion relations do not become identical, because of $\delta p_{\parallel}\neq\delta p_{\perp}$.